# Flips in colorful triangulations

### v1

@@ 6. Proof of Theorems 8 and 6
We represent a $k$-ary tree with vertices $1,\ldots,n$ by maintaining an array of length $k$ for the children of each vertex. Furthermore, each vertex has access to its parent vertex. With these data structures, up- and down-rotations by 1 step can be performed in time ${\mathcal{O}}(k)$. The algorithm to generate all trees by rotations is a faithful implementation of the inductive zigzag construction described in Section 5. As we are only computing a Hamilton path in the rotation graph (not a cycle), the interleaving of children sequences for the special trees $S_{n}$ and $S_{n}^{\prime}$ discussed in the proof of Theorem 7 is omitted, which simplifies this computation. Specifically, the selection of the vertex to be rotated next and in which direction can be achieved in time ${\mathcal{O}}(1)$ using simple data structures developed for the Hartung-Hoang-Mütze-Williams permutation language framework; see e.g. [MM23, CMM22].
@@ 8. Open questions
Another interesting question concerns bijections between $k$-ary trees and classes of permutations. For $k=2$ (binary trees), there is a natural bijection to 231-avoiding permutations. Are there similar correspondences between $k$-ary trees and pattern-avoiding permutations for $k\geq 3$? In particular, do tree rotations translate to nice operations on the permutations, specifically to so-called jumps heavily used in [HHMW22, HM21, MM23, CMM22]?

### v2

@@ 6. Proofs of Theorems 8 and 6
We represent a $k$-ary tree with vertices $1,\ldots,n$ by maintaining an array of length $k$ for the children of each vertex. Furthermore, each vertex has access to its parent vertex. With these data structures, up- and down-rotations by 1 step can be performed in time ${\mathcal{O}}(k)$. The algorithm to generate all trees by rotations is a faithful implementation of the inductive zigzag construction described in Section 5. As we are only computing a Hamilton path in the rotation graph (not a cycle), the interleaving of children sequences for the special trees $S_{n}$ and $S_{n}^{\prime}$ discussed in the proof of Theorem 7 is omitted, which simplifies this computation. Specifically, the selection of the vertex to be rotated next and in which direction can be achieved in time ${\mathcal{O}}(1)$ using simple data structures developed for the Hartung-Hoang-Mütze-Williams permutation language framework; see e.g. [MM23, CMM24].
@@ 8. Open questions
Another interesting question concerns bijections between $k$-ary trees and classes of permutations. For $k=2$ (binary trees), there is a natural bijection to 231-avoiding permutations. Are there similar correspondences between $k$-ary trees and pattern-avoiding permutations for $k\geq 3$? In particular, do tree rotations translate to nice operations on the permutations, specifically to so-called jumps heavily used in [HHMW22, HM21, MM23, CMM24]?
@@
Baur, Bergerova, Voon and Xu [BBVX24] recently introduced another family of flip graphs on triangulations in which the triangles are colored, not the vertices. The resulting graphs are disconnected in general, and their structure is still not very well understood (in [GP02] these graphs are related to the famous Four Color Theorem).